OpenStudy (anonymous):

Exponential decay, a paleontologist finds a bone that might be a dinosaur bone. in the laboratory she finds that the carbon-14 found in the bone is 1/12 of that found in living bone tissue. Could this bone have belonged to a dinosaur? Explain your reasoning. Hint- dinosaurs lived from 220 million years ago to 63 million years ago. Please show work!

OpenStudy (anonymous):

would help if you provided the half life of carbon 14 i think it is something like 5700 years, but i forget

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then we need to solve \[\frac{1}{12}=\left(\frac{1}{2}\right)^{\frac{t}{5700}}\] and see if t is in that interval. do you know how to solve it?

OpenStudy (anonymous):

yea i know how to solve it just didn't know the half life thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

You gave me the wrong formula.

OpenStudy (anonymous):

Its actually y=ae^-.0012t!

OpenStudy (anonymous):

i am confuse